OpenStudy (anonymous):

Below is Brianna’s daily schedule. For each activity, identify the elapsed time from its beginning to its end. Be sure to show your work. Finally, explain why your answer is correct. Schoolwork: 7:35 a.m. until 11:15 a.m.

OpenStudy (anonymous):

7:35 -10:35 = 180 min 10:35 - 11:00 = 25 min 11:00 - 11:15 = 15 min Total the number of minutes = 220 220 minutes/60 = 3h 40 min.

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

welcome sis